the equations are :
y=4x-2
y=x-2
what is the solution of the system of equations and is it consistent or not
I never learned this and i have to do an online course for keystone exams so i am really lost. Please help me

Answer by tommyt3rd(5050)

y=4x-2
y=x-2

4x-2=x-2
x=0
y=-2
system is consistent,
solution (0,-2)